OpenStudy (anonymous):

What polynomial has roots of -5, -4, and 1 ? x^3 - 8x^2 - 11x + 20 x^3 - x^2 + 22x + 40 x^3 + x^2 - 22x - 40 x^3 + 8x^2 + 11x - 20

OpenStudy (anonymous):

@KirbyLegs last question i promise lol

OpenStudy (anonymous):

@sourwing could you help with this?

OpenStudy (anonymous):

\[\left( x+5 \right)\left( x+4 \right)\left( x-1 \right)=0\] solve.

OpenStudy (anonymous):

solve means multiply

OpenStudy (anonymous):

im doing it now, wait 1 sec

OpenStudy (anonymous):

its D?

OpenStudy (anonymous):

yes Here 5*4*-1=-20 which is only D option If you have more options with -20 ,then you have to proceed as above.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw